OpenStudy (anonymous):

will award medal Find the linear approximation of the given function at (2π, 0). f(x,y)=sqrt(y + (cos(x))^2 @SolomonZelman

OpenStudy (solomonzelman):

don't get the function

OpenStudy (anonymous):

\[f(x,y)=\sqrt{y+(\cos(x))^2}\]

OpenStudy (solomonzelman):

then you left out a bracket. but good that you fixed it.

OpenStudy (anonymous):

so my partial with respect to x i think is \[\frac{ 2\cos(x)(-\sin(x)) }{ 2\sqrt{y+(\cos(x))^2} }\]

OpenStudy (solomonzelman):

looks fine, the y on top those is 0?

OpenStudy (solomonzelman):

derivative of y is 1, no ?

OpenStudy (solomonzelman):

the chain for cos^2(x) is good

OpenStudy (anonymous):

no because the derivitive of y with respect to x is zero, because y is a constant

OpenStudy (anonymous):

or treated as one

OpenStudy (solomonzelman):

oh, ttrruuee

OpenStudy (solomonzelman):

yes, it is just not something I used to though:)

OpenStudy (solomonzelman):

Yes.

OpenStudy (anonymous):

but for my derivitive with respect to y, would you use the chain rule for (cos(x))^2, and instead of f(x)=x^2 and g(x)=cos(x)... f(y)=y^2, g(y)=cos(x)?

OpenStudy (solomonzelman):

if you are differentiating with respect to y, then the cos(x) is constant.

OpenStudy (solomonzelman):

and so is cos^2x

OpenStudy (anonymous):

so f_y would just be 1?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

the square root has a variable inside (the y), so differentiate the root, and apply chain

OpenStudy (anonymous):

it would be sqrtx

OpenStudy (solomonzelman):

I have to go right now to hear a guest speaker in my community. Excuse me

jaynator495 (jaynator495):

im not stalking anyone... you have no proof

OpenStudy (solomonzelman):

1/(sqrt of this) times derivative of what is in the root.

OpenStudy (solomonzelman):

I mean 1/(2( sqrt of that ))

OpenStudy (solomonzelman):

i gogt to go... I will be back in an hour.... sorry again.

OpenStudy (anonymous):

\[f(x,y)=\sqrt{y+\cos^2x}~~\implies~~\nabla f(x,y)=\left(-\frac{\cos x\sin x}{\sqrt{y+\cos^2x}},\,\frac{1}{2\sqrt{y+\cos^2x}}\right)\] At \((2\pi,0)\), you have \(\nabla f=\left(0,\dfrac{1}{2}\right)\), and so the linear approximation is given by \[z-f(2\pi,0)=0(x-2\pi)+\frac{1}{2}(y-0)\]

OpenStudy (anonymous):

Let me know if that code is showing up again.

OpenStudy (anonymous):

yea it was

OpenStudy (anonymous):

could you do the partial with respect to y step by step?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[f(x,y)=\sqrt{y+\cos^2x}\] \[f_y=\frac{1}{2}(y+\cos^2x)^{-1/2}\frac{\partial }{\partial y}[y+\cos^2x]=\frac{1}{2\sqrt{y+\cos^2x}}\]

OpenStudy (anonymous):

That is, power rule and chain rule.